OpenStudy (anonymous):

If a child's knowledge of the alphabet is limited to the letters a, b, c, i, and e, and if the child writes two letters at random (assume the child may write the same letter twice), what is the probability that one is a vowel and the other is a consonant? 6/25 12/25 4/5

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

4/5?

ganeshie8 (ganeshie8):

how ?

OpenStudy (anonymous):

well cause theres 5 letters

OpenStudy (anonymous):

wait its actaully 12/25

ganeshie8 (ganeshie8):

simply find the probabilities of each event and multiply

ganeshie8 (ganeshie8):

what is the probability that `one is a vowel` and the `other is a consonant`?

ganeshie8 (ganeshie8):

how many vowels are there in the given set of letters ?

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

got it thnx

ganeshie8 (ganeshie8):

you seem to be in rush lol

OpenStudy (anonymous):

i am, got a lot to do till tomorrow

ganeshie8 (ganeshie8):

i wish you all the best, this is an easy question : there are 3 vowels in a total of 5 letters so probability for vowel would be 3/5 similarly probabiltiy of consonant would be 2/5 multiplying them gives you the desired probability of having a vowel and consonant : (3/5)*(2/5) = 6/25

OpenStudy (anonymous):

it isn't 6/25

ganeshie8 (ganeshie8):

the grader marked it wrong is it ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

Oh right, silly mistake! we can have the first letter as "consonant" so the probability would be \[(3/5)*(2/5) + (2/5)*(3/5) = 12/25\]

OpenStudy (anonymous):

oh ok thnx

ganeshie8 (ganeshie8):

yw!

OpenStudy (freckles):

if you don't like that you could have also wrote it all out: